OpenStudy (anonymous):

anyone want to help me please? What are the ordered pairs? Using the process of elimination method. 5u +3v = -18 2u + v = -8

OpenStudy (anonymous):

multiply the second equation by 3 all the way across and subtract it from the first equation

OpenStudy (anonymous):

or if you prefer multiply the second equation by \(-3\) and add. here are the steps

OpenStudy (anonymous):

\[5u+3v=-18\] \[-6u-3v=24\] now add in a column, get \[-u=6\] and so \(u=-6\)

OpenStudy (anonymous):

I get that part but I am confused on how to find the ordered pairs?

OpenStudy (anonymous):

oh now that you have \(u=-6\) you can find \(v\) by substituting \(-6\) for \(u\) in either equation

OpenStudy (anonymous):

\[2u + v = -8 \] \[2(-6)+v=-8\] \[-12+v=-8\] \[v=4\]

OpenStudy (anonymous):

so my ordered pairs are (-6,4) ?

OpenStudy (anonymous):

so the "ordered pair" is \((-6,4)\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you i am glad you showed the ''steps'' as well. so that I understand. :)

OpenStudy (anonymous):

yw